# Differentially-Private Distributed Model Predictive Control of Linear Discrete-Time Systems with Global Constraints

### v1

@@ 3.3 Privacy Analysis
For Algorithm 3, an execution is represented as $\mathcal{A}=\{\vartheta^{0},\vartheta^{1},\ldots\}$ with $\vartheta^{k}=\lambda^{k}=\begin{bmatrix}(\lambda_{1}^{k})^{\top},\cdots,(\lambda_{M}^{k})^{\top}\end{bmatrix}^{\top}$. An observation sequence is denoted as $\mathcal{O}=\{o^{0},o^{1},\ldots\}$ with $o^{k}=\hat{\lambda}^{k}=\begin{bmatrix}(\hat{\lambda}_{1}^{k})^{\top},\cdots,(\hat{\lambda}_{M}^{k})^{\top}\end{bmatrix}^{\top}$ (note that $\hat{\lambda}_{i}^{k}=\lambda_{i}^{k}+\zeta_{i}^{k}$, as detailed in Algorithm 3). Similar to the sensitivity definition of constraint-free distributed optimization in Huang et al. (2015), we define the sensitivity of an algorithm as follows:
@@
At each iteration $k$, for any initial state $\vartheta^{0}$ and adjacent distributed optimization problems $\mathcal{P}$ and $\mathcal{P}^{{}^{\prime}}$, the sensitivity of Algorithm 3 is
@@
The proof of the lemma follows the same line of reasoning as that of Lemma 2 in Huang et al. (2015).
We also need the following lemma for privacy analysis:
@@
(Lemma 4, Chung (1954)) Let $\{\psi^{k}\}$ be a non-negative sequence, and $\{a^{k}\}$ and $\{b^{k}\}$ be positive sequences satisfying $\sum_{k=0}^{\infty}a^{k}=\infty$, $\lim_{k\rightarrow\infty}a^{k}=0$, and $\frac{b^{k}}{a^{k}}$ converges to zero with a polynomial rate. If there exists a $K\geq 0$ such that $\psi^{k+1}\leq(1-a^{k})\psi^{k}+b^{k}$ holds for all $k\geq K$, then we always have $\psi^{k}\leq\bar{C}\frac{b^{k}}{a^{k}}$ for all $k$, where $\bar{C}$ is some constant.
@@
Suppose the conditions of Theorem 1 hold. If all elements of $\zeta_{i}^{k}$ are drawn independently from Laplace distribution ${\rm Lap}(\nu^{k})$ with $(\sigma_{i}^{k})^{2}=2(\nu^{k})^{2}$ satisfying Assumption 4, then the following results hold:
@@
    For any finite number of iterations $T$, Algorithm 3 is $\epsilon$-differentially private with the cumulative privacy budget bounded by $\epsilon\leq\sum_{k=1}^{T}\frac{C\varsigma^{k}}{\nu^{k}}$, where $\varsigma^{k}:=\sum_{s=1}^{k-1}\Pi_{q=s}^{k-1}(1-\chi^{q}\bar{L})\gamma^{s-1}\chi^{s-1}+\gamma^{k-1}\chi^{k-1}$, $\bar{L}:=min_{i\in\mathbb{Z}_{1}^{M}}|L_{ii}|$, and $C$ is from (34);
@@
    If the sequence $\{\frac{\gamma^{k}}{\nu^{k}}\}$ is summable, then the cumulative privacy budget is finite as $T\rightarrow\infty$.
@@
To establish the privacy guarantees, we begin by analyzing the sensitivity of Algorithm 3. Given two adjacent distributed optimization problems $\mathcal{P}$ and $\mathcal{P}^{{}^{\prime}}$, and considering a fixed observation $\mathcal{O}$ and initial state $\lambda^{0}$, the sensitivity depends on $\|\lambda^{k}-\lambda^{\prime k}\|_{1}$ as per Definition 3. Since $\mathcal{P}$ and $\mathcal{P}^{{}^{\prime}}$ differ only in one signal, we denote this distinct signal as the $i$th one, i.e., $g_{i}$ in $\mathcal{P}$ and $g^{{}^{\prime}}_{i}$ in $\mathcal{P}^{{}^{\prime}}$, without loss of generality. Since the initial conditions and observations of $\mathcal{P}$ and $\mathcal{P}^{{}^{\prime}}$ are identical for $j\neq i$, we have $\lambda_{j}^{k}=\lambda^{\prime k}_{j}$ for all $j\neq i$ and $k$, which indicates that $\|\lambda^{k}-\lambda^{\prime k}\|_{1}$ is always equal to $\|\lambda_{i}^{k}-\lambda^{\prime k}_{i}\|_{1}$.
Based on (18) in Algorithm 3, $L_{ii}:=-\sum_{j\in\mathbb{N}_{i}}L_{ij}$, and the fact that the observations $\lambda_{j}^{k}+\zeta_{j}^{k}$ and ${\lambda^{\prime}_{j}}^{k}+{\zeta^{\prime}_{j}}^{k}$ are the same, we can derive that
$\|\lambda_{i}^{k+1}-{\lambda^{\prime}_{i}}^{k+1}\|_{1}\leq(1-|L_{ii}|\chi^{k})\|\lambda_{i}^{k}-{\lambda^{\prime}_{i}}^{k}\|_{1}+\gamma^{k}\|g_{i}(\tilde{\bm{u}}_{i}^{k+1})-{g^{{}^{\prime}}_{i}(\tilde{\bm{u}}^{\prime k+1}_{i})}\|_{1}.$ (36)
Therefore, it can be obtained from (34) and (36) that the sensitivity $\Delta^{k}$ satisfies
@@
$\leq(1-|L_{ii}|\chi^{k})\Delta^{k}+C\gamma^{k}\chi^{k},$
which implies the first statement by iteration using Lemma 2.
Lemma 3 is exploited to prove the infinity horizon result in the second statement. Specifically, based on (37) and the properties of $\chi^{k}$ and $\gamma^{k}$, Lemma 3 can be used to conclude that there exists some $\bar{C}$ such that $\Delta^{k}$ satisfies $\Delta^{k}\leq\bar{C}\gamma^{k}$. It can be further obtained from Lemma 2 that $\epsilon\leq\sum_{k=1}^{T}\frac{\bar{C}\gamma^{k}}{\nu^{k}}$. Thus, $\epsilon$ will be finite even when $T$ tends to infinity if $\sum_{k=0}^{\infty}\frac{\gamma^{k}}{\nu^{k}}<\infty$ holds.
4 Implementation of Privacy-Preserving DMPC
@@ 4.1 Algorithm Implementation
To verify whether the global constraints are satisfied after the termination of Algorithm 3, we employ a privacy-preserving static average consensus method developed in Wang (2019).
@@
where $\iota$, $a_{i,\alpha\beta}^{\ell}$, $a_{i,j}^{\ell}\in\mathbb{R}_{+}$. As proven in Wang (2019), by appropriately selecting the parameters $\iota$, $a_{i,\alpha\beta}^{\ell}$, and $a_{i,j}^{\ell}$, $z_{i,\alpha}^{\ell}$ and $z_{i,\beta}^{\ell}$ converge to the average consensus value $\frac{1}{M}\sum_{i=1}^{M}z_{i}^{0}$ (i.e., $\frac{1}{M}\sum_{i=1}^{M}g_{i}(\tilde{{\bm{u}}}_{i}^{\bar{k}})$). Therefore, each subsystem can utilize the converged value of $z_{i,\alpha}^{\ell}$ to check whether condition (38) is satisfied. It is worth noting that conventional static average consensus approaches (Olfati-Saber et al., 2007; Sundaram and Hadjicostis, 2007; Hendrickx et al., 2015) can also be employed to calculate the value of $\frac{1}{M}\sum_{i=1}^{M}z_{i}^{0}$ in a distributed manner. However, these approaches necessitate subsystems to directly share $z_{i}^{0}$ with their neighbors, potentially leading to privacy breaches as $z_{i}^{0}=g_{i}(\tilde{{\bm{u}}}_{i}^{\bar{k}})$ contains sensitive information about $\tilde{{\bm{u}}}_{i}^{\bar{k}}$. The average consensus method developed in Wang (2019) employs a state decomposition scheme to mask the real values of $z_{i}^{0}$. As shown in (39), the substate $z_{i,\alpha}^{\ell}$ governs the role of internode interactions and is the only value from subsystem $i$ that can be seen by its neighbors. On the other hand, the other substate $z_{i,\beta}^{\ell}$ participates in the distributed interactions by solely interacting with $z_{i,\alpha}^{\ell}$. Hence, the existence of $z_{i,\beta}^{\ell}$ is invisible to neighboring nodes of subsystem $i$, although it directly affects the evolution of $z_{i,\alpha}^{\ell}$. Through this state decomposition design, strong privacy preservation can be guaranteed. For further details, please refer to Wang (2019).
@@
1 At time instant $t$, each subsystem $i$ measures its state $x_{i}(t)$;
2Each subsystem $i$ obtains $\tilde{{\bm{u}}}_{i}^{\bar{k}}$ by following Algorithm 3 with $x_{i}(t)$;
3Each subsystem $i$ runs the static average consensus algorithm (39) to obtain $\sum_{i=1}^{M}g_{i}(\tilde{{\bm{u}}}_{i}^{\bar{k}})$;
@@ 1 At time instant $t$, each subsystem $i$ measures its state $x_{i}(t)$;
Wait for the next sampling instant; let $t=t+1$ and go to step 1.
@@ 4.2 Feasibility and Stability
    If Algorithm 4 has a feasible solution at time $t$, then it has a feasible solution at $t+1$.
@@
As shown in Algorithm 4, the input sequence at time instant $t$ is denoted by $\tilde{{\bm{u}}}_{i}(t)=\{\tilde{u}_{i}(0|t),\tilde{u}_{i}(1|t),\cdots,\tilde{u}_{i}(N-1|t)\}$. The corresponding predicted state sequence is $\tilde{{\bm{x}}}_{i}(t)=\{\tilde{x}_{i}(0|t),\tilde{x}_{i}(1|t),\cdots,\tilde{x}_{i}(N|t)\}$. Since $\tilde{{\bm{u}}}_{i}(t)$ is a feasible solution, it can be obtained from (7), (8), and (38) that $\tilde{{\bm{u}}}_{i}(t)\in\tilde{\mathcal{U}_{i}}(x_{i}(t))$ and
@@
At time instant $t+1$, an input sequence $\hat{{\bm{u}}}_{i}(t+1)$ and its corresponding predicted state sequence $\hat{{\bm{x}}}_{i}(t+1)$ are defined as follows:
$\hat{{\bm{u}}}_{i}(t+1)$ =$=\{\hat{u}_{i}(0|t+1),\hat{u}_{i}(1|t+1),\cdots,\hat{u}_{i}(N-1|t+1)\}$ (42)
=$=\{\tilde{u}_{i}(1|t),\tilde{u}_{i}(2|t),\cdots,\tilde{u}_{i}(N-1|t),K_{i}\tilde{x}_{i}(N|t)\},$
$\hat{{\bm{x}}}_{i}(t+1)$ =$=\{\hat{x}_{i}(0|t+1),\hat{x}_{i}(1|t+1),\cdots,\hat{x}_{i}(N|t+1)\}$
=$=\{\tilde{x}_{i}(1|t),\tilde{x}_{i}(2|t),\cdots,\tilde{x}_{i}(N|t),(A_{i}+B_{i}K_{i})\tilde{x}_{i}(N|t)\}.$
@@
$J_{i}(x_{i}(t+1),\hat{{\bm{u}}}_{i}(t+1))-J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i}(t))=-\|x_{i}(t)\|^{2}_{Q_{i}}-\|\tilde{u}_{i}(0|t)\|^{2}_{R_{i}}.$ (44)
As the feasible solution $\hat{{\bm{u}}}_{i}(t+1)$ may not be optimal at $t+1$,
$V(\{x_{i}(t+1)\})$ $\leq\sum_{i=1}^{M}J_{i}(x_{i}(t+1),\hat{{\bm{u}}}_{i}(t+1))$ (45)
@@
where the equality condition is due to (44) and the last inequality follows from (40). (45) indicates that $x_{i}(t)$ converges to the bounded set $\{\{x_{i}\}:\sum_{i=1}^{M}\|x_{i}\|^{2}_{Q_{i}}\leq\eta\}$ in finite time. Considering this fact and the assumption that $\{x_{i}\in\mathbb{R}^{n_{i}}:\|x_{i}\|^{2}_{Q_{i}}\leq\eta\}\subset\mathcal{X}^{f}_{i}$, it can be concluded that $x_{i}(t)$ enters the terminal set $x_{i}(t)$ in finite time.
5 Numerical Simulations
In this section, simulation is conducted to demonstrate the performance of the developed approach. A group of four linear time-invariant subsystems are considered. The network structure of these four subsystems is shown in Figure 1. The system matrices $A_{i}$ and $B_{i}$ are chosen as

### v2

@@ III-C Privacy Analysis
For Algorithm 3, an execution is represented as $\mathcal{A}=\{\vartheta^{0},\vartheta^{1},\ldots\}$ with $\vartheta^{k}=\lambda^{k}=\begin{bmatrix}(\lambda_{1}^{k})^{\top},\cdots,(\lambda_{M}^{k})^{\top}\end{bmatrix}^{\top}$. An observation sequence is denoted as $\mathcal{O}=\{o^{0},o^{1},\ldots\}$ with $o^{k}=\hat{\lambda}^{k}=\begin{bmatrix}(\hat{\lambda}_{1}^{k})^{\top},\cdots,(\hat{\lambda}_{M}^{k})^{\top}\end{bmatrix}^{\top}$ (note that $\hat{\lambda}_{i}^{k}=\lambda_{i}^{k}+\zeta_{i}^{k}$, as detailed in Algorithm 3). Similar to the sensitivity metric proposed for constraint-free distributed optimization in [22], we formulate the sensitivity of Algorithm 3 in the following manner:
@@
At each iteration $k$, for any two adjacent distributed optimization problems $\mathcal{P}$ and $\mathcal{P}^{{}^{\prime}}$ and any initial state $\vartheta^{0}$, the sensitivity of Algorithm 3 is given by
@@
The proof of this lemma follows the same reasoning as that of Lemma 2 in [22]. ∎
We also introduce the following lemma for privacy analysis:
@@
(Lemma 4, [33]) Let $\{\psi^{k}\}$ be a non-negative sequence, and $\{a^{k}\}$ and $\{b^{k}\}$ be positive sequences satisfying $\sum_{k=0}^{\infty}a^{k}=\infty$, $\lim_{k\rightarrow\infty}a^{k}=0$, and $\frac{b^{k}}{a^{k}}$ converges to zero with a polynomial rate. If there exists a $\bar{K}\geq 0$ such that $\psi^{k+1}\leq(1-a^{k})\psi^{k}+b^{k}$ holds for all $k\geq\bar{K}$, then it follows that $\psi^{k}\leq\bar{C}\frac{b^{k}}{a^{k}}$ for all $k$, with $\bar{C}$ being some constant.
@@
Suppose the conditions of Theorem 1 hold. If every element of $\zeta_{i}^{k}$ is independently sampled from Laplace distribution ${\rm Lap}(\nu^{k})$, where $(\sigma_{i}^{k})^{2}=2(\nu^{k})^{2}$ satisfies Assumption 4, then the following results hold:
@@
    For any finite number of iterations $T$, Algorithm 3 ensures $\epsilon$-differential privacy, and the cumulative privacy budget is bounded by $\epsilon\leq\sum_{k=1}^{T}\frac{C\varsigma^{k}}{\nu^{k}}$, where $\varsigma^{k}:=\sum_{s=1}^{k-1}\Pi_{q=s}^{k-1}(1-\chi^{q}\bar{L})\gamma^{s-1}\chi^{s-1}+\gamma^{k-1}\chi^{k-1}$, $\bar{L}:=\min_{i\in\mathbb{Z}_{1}^{M}}|L_{ii}|$, and $C$ is from (34);
@@
    If $\sum_{k=0}^{\infty}\frac{\gamma^{k}}{\nu^{k}}<\infty$ holds, then the cumulative privacy budget remains finite as $T\rightarrow\infty$.
@@
To establish the privacy guarantees, we begin by analyzing the sensitivity of Algorithm 3. Given any initial state $\lambda^{0}$, any fixed observation $\mathcal{O}$, and two adjacent distributed optimization problems $\mathcal{P}$ and $\mathcal{P}^{{}^{\prime}}$, the sensitivity depends on $\|\lambda^{k}-\lambda^{\prime k}\|_{1}$ as per Definition 3. Note that $\mathcal{P}$ and $\mathcal{P}^{{}^{\prime}}$ differ solely in one signal, and without loss of generality, we denote this distinct signal as the $i$th one, i.e., $g_{i}$ in $\mathcal{P}$ and $g^{{}^{\prime}}_{i}$ in $\mathcal{P}^{{}^{\prime}}$. Since the initial conditions and observations of $\mathcal{P}$ and $\mathcal{P}^{{}^{\prime}}$ are the same for $j\neq i$, it follows that $\lambda_{j}^{k}=\lambda^{\prime k}_{j}$ for all $k$ and $j\neq i$. Consequently, $\|\lambda^{k}-\lambda^{\prime k}\|_{1}$ is always equal to $\|\lambda_{i}^{k}-\lambda^{\prime k}_{i}\|_{1}$.
Based on (18) in Algorithm 3, $L_{ii}:=-\sum_{j\in\mathbb{N}_{i}}L_{ij}$, and the fact that the observations $\lambda_{j}^{k}+\zeta_{j}^{k}$ and ${\lambda^{\prime}_{j}}^{k}+{\zeta^{\prime}_{j}}^{k}$ are identical, we can derive that
$\|\lambda_{i}^{k+1}-{\lambda^{\prime}_{i}}^{k+1}\|_{1}\leq$ $(1-|L_{ii}|\chi^{k})\|\lambda_{i}^{k}-{\lambda^{\prime}_{i}}^{k}\|_{1}$ (36)
$\quad+\gamma^{k}\|g_{i}(\tilde{\bm{u}}_{i}^{k+1})-{g^{{}^{\prime}}_{i}(\tilde{\bm{u}}^{\prime k+1}_{i})}\|_{1}.$
Therefore, it can be obtained from (34) and (36) that the sensitivity $\Delta^{k}$ is bounded by
@@
$\leq(1-|L_{ii}|\chi^{k})\Delta^{k}+C\gamma^{k}\chi^{k}.$
According to Lemma 2 and (37), the first statement can be obtained.
Lemma 3 is exploited to prove the second statement of Theorem 3. Specifically, based on (37) and the properties of $\chi^{k}$ and $\gamma^{k}$, Lemma 3 can be used to conclude that there exists some constant $\bar{C}$ such that the sensitivity $\Delta^{k}$ satisfies $\Delta^{k}\leq\bar{C}\gamma^{k}$. It can be further obtained from Lemma 2 that $\epsilon\leq\sum_{k=1}^{T}\frac{\bar{C}\gamma^{k}}{\nu^{k}}$. Thus, if $\sum_{k=0}^{\infty}\frac{\gamma^{k}}{\nu^{k}}<\infty$ holds (i.e., the sequence $\{\frac{\gamma^{k}}{\nu^{k}}\}$ is summable), then $\epsilon$ will be finite even when $T\rightarrow\infty$. ∎
IV Implementation of Privacy-Preserving DMPC
@@ IV-A Algorithm Implementation
To verify whether the global constraints are satisfied after the termination of Algorithm 3, we employ a privacy-preserving static average consensus method developed in [34].
@@
where $\iota$, $a_{i,\alpha\beta}^{\ell}$, $a_{i,j}^{\ell}\in\mathbb{R}_{+}$. As proven in [34], by appropriately selecting the parameters $\iota$, $a_{i,\alpha\beta}^{\ell}$, and $a_{i,j}^{\ell}$, $z_{i,\alpha}^{\ell}$ and $z_{i,\beta}^{\ell}$ converge to the average consensus value $\frac{1}{M}\sum_{i=1}^{M}z_{i}^{0}$ (i.e., $\frac{1}{M}\sum_{i=1}^{M}g_{i}(\tilde{{\bm{u}}}_{i}^{\bar{k}})$). Therefore, each subsystem can utilize the converged value of $z_{i,\alpha}^{\ell}$ to check whether condition (38) is satisfied. It is worth noting that conventional static average consensus approaches [35, 36, 37] can also be employed to calculate the value of $\frac{1}{M}\sum_{i=1}^{M}z_{i}^{0}$ in a distributed manner. However, these approaches necessitate subsystems to directly share $z_{i}^{0}$ with their neighbors, potentially leading to privacy breaches as $z_{i}^{0}=g_{i}(\tilde{{\bm{u}}}_{i}^{\bar{k}})$ contains sensitive information about $\tilde{{\bm{u}}}_{i}^{\bar{k}}$. The average consensus method developed in [34] employs a state decomposition scheme to mask the real values of $z_{i}^{0}$. As shown in (39), the substate $z_{i,\alpha}^{\ell}$ governs the role of internode interactions and is the only value from subsystem $i$ that can be seen by its neighbors. On the other hand, the other substate $z_{i,\beta}^{\ell}$ participates in the distributed interactions by solely interacting with $z_{i,\alpha}^{\ell}$. Hence, the existence of $z_{i,\beta}^{\ell}$ is invisible to neighboring nodes of subsystem $i$, although it directly affects the evolution of $z_{i,\alpha}^{\ell}$. Through this state decomposition design, strong privacy preservation can be guaranteed. For further details, please refer to [34].
@@
1 At time instant $t$, every subsystem $i$ measures its state $x_{i}(t)$;
2Every subsystem $i$ computes $\tilde{{\bm{u}}}_{i}^{\bar{k}}$ by following Algorithm 3 with $x_{i}(t)$;
3Every subsystem $i$ runs the static average consensus algorithm (39) to obtain $\sum_{i=1}^{M}g_{i}(\tilde{{\bm{u}}}_{i}^{\bar{k}})$;
@@ 1 At time instant $t$, every subsystem $i$ measures its state $x_{i}(t)$;
Wait for the next time instant; let $t=t+1$ and go to step 1.
@@ IV-B Feasibility and Stability
    If Algorithm 4 has a feasible solution at time instant $t$, then it has a feasible solution at $t+1$.
@@
As shown in Algorithm 4, the input sequence at time instant $t$ is denoted by $\tilde{{\bm{u}}}_{i}(t)=\{\tilde{u}_{i}(0|t),\tilde{u}_{i}(1|t),\cdots,\tilde{u}_{i}(N-1|t)\}$. Let $\tilde{{\bm{x}}}_{i}(t)=\{\tilde{x}_{i}(0|t),\tilde{x}_{i}(1|t),\cdots,\tilde{x}_{i}(N|t)\}$ be the corresponding predicted state sequence. Since $\tilde{{\bm{u}}}_{i}(t)$ is a feasible solution, it can be obtained from (7), (8), and (38) that $\tilde{{\bm{u}}}_{i}(t)\in\tilde{\mathcal{U}_{i}}(x_{i}(t))$ and
@@
At time instant $t+1$, an input sequence $\hat{{\bm{u}}}_{i}(t+1)$ and its corresponding predicted state sequence $\hat{{\bm{x}}}_{i}(t+1)$ are defined as
$\hat{{\bm{u}}}_{i}(t+1)$ (42)
$=$ $\{\hat{u}_{i}(0|t+1),\hat{u}_{i}(1|t+1),\cdots,\hat{u}_{i}(N-1|t+1)\}$
$=$ $\{\tilde{u}_{i}(1|t),\tilde{u}_{i}(2|t),\cdots,\tilde{u}_{i}(N-1|t),K_{i}\tilde{x}_{i}(N|t)\},$
$\hat{{\bm{x}}}_{i}(t+1)$
$=$ $\{\hat{x}_{i}(0|t+1),\hat{x}_{i}(1|t+1),\cdots,\hat{x}_{i}(N|t+1)\}$
$=$ $\{\tilde{x}_{i}(1|t),\tilde{x}_{i}(2|t),\cdots,\tilde{x}_{i}(N|t),(A_{i}+B_{i}K_{i})\tilde{x}_{i}(N|t)\}.$
@@
$J_{i}(x_{i}(t+1),\hat{{\bm{u}}}_{i}(t+1))-J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i}(t))$ (44)
$=$ $-\|x_{i}(t)\|^{2}_{Q_{i}}-\|\tilde{u}_{i}(0|t)\|^{2}_{R_{i}}.$
$\hat{{\bm{u}}}_{i}(t+1)$ is a feasible solution at $t+1$ but may not be optimal. Thus, we have
$V(\{x_{i}(t+1)\})$ (45)
$\leq$ $\sum_{i=1}^{M}J_{i}(x_{i}(t+1),\hat{{\bm{u}}}_{i}(t+1))$
@@
where the equality condition is due to (44) and the last inequality follows from (40). (45) indicates that $x_{i}(t)$ converges to the bounded set $\{\{x_{i}\}:\sum_{i=1}^{M}\|x_{i}\|^{2}_{Q_{i}}\leq\eta\}$ in finite time. Considering this fact and the assumption that $\{x_{i}\in\mathbb{R}^{n_{i}}:\|x_{i}\|^{2}_{Q_{i}}\leq\eta\}\subset\mathcal{X}^{f}_{i}$, it can be concluded that $x_{i}(t)$ enters the terminal set $x_{i}(t)$ in finite time. ∎
V Numerical Simulations
In this section, simulation is conducted to demonstrate the performance of the developed method. A group of four linear time-invariant subsystems are considered. The network structure of these four subsystems is shown in Figure 1. The system matrices $A_{i}$ and $B_{i}$ are chosen as